OpenStudy (anonymous):

The height of a helicopter above the ground is given by h = 2.80t^3, where h is in meters and t is in seconds. At t = 2.50 s, the helicopter releases a small mailbag. How long after its release does the mailbag reach the ground?

OpenStudy (shane_b):

First, you need to calculate the height when it gets dropped. Can you do that part?

OpenStudy (anonymous):

yeah, height should be 43.75

OpenStudy (shane_b):

Right. So from there you can just use one of the kinematic equations:\[d=V_ot+\frac{1}{2}gt^2\]Here, V0 is 0 m/s since it's being dropped. So you end up with this equation:\[43.75m=\frac{1}{2}(9.8m/s^2)t^2\]Now you just need to solve it for t.

OpenStudy (anonymous):

t should be t=2.988, but webassign says its wrong... Apparently the correct answer is 11.5

OpenStudy (shane_b):

With the information given, there's no way it would 11.5 seconds. That doesn't even make logical sense.

OpenStudy (anonymous):

then ask webassign, cause it's saying 2.988 is wrong

OpenStudy (shane_b):

Maybe there's a typo in the question...either by you or in webassign. That's all I can think of. This is a very basic physics question.

OpenStudy (anonymous):

i thought so too until i couldnt even get the "right answer"

OpenStudy (shane_b):

Wait...I think this is supposed to be a projectile motion question. That means you also have to account for the upward velocity of the helicopter in that it reached a height of 43.75m in 2.0s.

OpenStudy (anonymous):

sadly, i still dont know what to do from there

OpenStudy (shane_b):

Gimme a minute...I'll do it that way. I made the incorrect assumption that the helicopter was hovering at the point at which the mailbag was dropped.

OpenStudy (shane_b):

\[H_0=2.80t^3=43.75m\]\[t=2.50s\]\[V_{0} = 8.4t^2=8.4(2.80)^2=52.5m/s\]The max height will be : \[H_{max}=H_0+V_0t+\frac{1}{2}gt^2\]\[0 = 43.75m + 52.5m/s + \frac{1}{2}(-9.8m/s^2)t^2\]Solving for t you end up with the times where it's at max height as well as when it hits the ground.

OpenStudy (shane_b):

You still there?

OpenStudy (anonymous):

yeah..

OpenStudy (shane_b):

Sorry...I had the correct value the first time. I'm just tired... \[0 = 43.75m + 52.5m/s(t) + \frac{1}{2}(-9.8m/s^2)t^2\]\[t=11.5s\]

OpenStudy (shane_b):

Ok...now that I'm done screwing up...do you understand what I did?

OpenStudy (anonymous):

wish i did.. haha

OpenStudy (anonymous):

i see that you calculated the velocity when the bag hit the ground using the derivative, etc but i dont see how you pulled 11.5 out of that equation

OpenStudy (shane_b):

Ok, we know the starting height is 43.75m. Since the helicopter has an upward velocity when the mailbag is dropped you need to calculate what that is. Take the derivative of 2.80t^3 to get 8.4t^2. Plugging in 2.5s as t gets you 52.5m/s. The rest is just using the kinematic equation:\[\large h_{final}=h_{initial}+V_0t+\frac{1}{2}gt^2\]

OpenStudy (anonymous):

using that equation is still get 5.976

OpenStudy (shane_b):

http://www.wolframalpha.com/input/?i=solve+for+t%3A+0+%3D+43.75+%2B+52.5t+%2B+ \frac{1}{2}%28-9.8%29t^2

OpenStudy (shane_b):

Copy the entire link...not sure why it didn't paste correctly.

OpenStudy (anonymous):

okay ill work on it, thanks

OpenStudy (shane_b):

At this point you just need to solve for t correctly using my last equation. You should get 11.491 which is about 11.5s.

OpenStudy (shane_b):

np